# Learning to Solve and Verify: A Self-Play Framework for Code and Test Generation

### v2

@@ 1 Introduction
Large language models (LLMs) have demonstrated impressive ability in code generation, significantly enhancing the programming efficiency and productivity of human developers (Li et al., 2022; Roziere et al., 2023; Chaudhary, 2023). The ability to code is largely driven by high-quality online resources where coding problems, human-written solutions, and corresponding unit tests are freely available. However, as these free resources being depleted, the momentum of LLMs’ improvement diminishes.
@@
To address the scarcity of supervised data for code generation, recent studies have adopted synthetic data generation techniques such as Self-Instruct (Wang et al., 2023) to augmenting LLM training sets. Specifically, previous work collects and designs code instructions and generates corresponding responses using a high-capacity teacher LLM. This generated data is then employed to fine-tune a student LLM, thereby enhancing its code generation abilities. Although synthetic code data produced in this manner has demonstrated success, it relies on the availability of a strong teacher model, presumably with a larger parameter size and higher computation costs. Additionally, existing work has shown that training a model on data generated by itself is ineffective because errors introduced during generation tend to accumulate over iterations (Dubey et al., 2024). As a result, there is a critical need for effective methods to verify the generated data.
Despite this pressing need, evaluating the correctness of the generated code is not trivial and often demands substantial programming expertise, even for human annotators. Recently, some research has explored the use of LLM-as-a-judge (McAleese et al., 2024; Alshahwan et al., 2024; Dong et al., 2024), which can automatically provide feedback on the code it has generated, for example by running the code against unit tests it produces. However, as we will show in this work, the current capability of LLMs to generate unit tests, herein referred to as LLM-as-a-verifier, is substantially worse than their ability to produce code solutions, herein referred to as LLM-as-a-solver (Section 4.3). This is because of the lack of high-quality data specifically for unit test generation during the post-training phase, i.e., most fine-tuning data focuses on code generation, while only a small portion targets unit test generation.
@@
In this work, we propose Sol-Ver, a self-play solver-verifier framework to iteratively train a model for both code and test generation. The main idea is to let the LLM-as-a-solver and LLM-as-a-verifier help each other. Specifically, we ask the model to generate code solutions and unit tests for the same set of coding problems. By executing the generated test against the generated code, we obtain feedback for training, involving two steps: (1) SFT training: we take the passed examples for fine-tuning the model, and (2) DPO training: we take both passed and failed examples as preference pairs to further train the model aligning with the preference. These training steps are for both code generation and unit test generation, and they can be repeated in an iterative manner.
@@
  • Identification of a critical gap: We analyze and highlight the significant gap in LLMs’ abilities between code generation and unit test generation.
@@ 2 Related Work
Various empirical studies applying scaling laws to the training of foundation models have highlighted the critical role of the data size (Kaplan et al., 2020; Hoffmann et al., 2022). To address the need for larger datasets, synthetic data generation has become a popular and cost-effective solution, which leverages advanced LLMs to produce high-quality data. One notable method is Self-Instruct (Wang et al., 2023), which employs a pre-trained LLM to generate instruction-output pairs from a small seed dataset.
@@
In the realm of code, previous work commonly devises synthetic coding instructions using a stronger teacher model (e.g., ChatGPT or GPT-4) and then finetunes a weaker student model (e.g., CodeAlpaca (Chaudhary, 2023) and CodeLlama (Roziere et al., 2023)) with the generated data to distill knowledge from the teacher. For example, code alpaca consists of 20K automatically generated code instructions by applying Self-Instruct on ChatGPT using 21 seed tasks. To further enhance the code abilities of LLMs, Luo et al. (2023) proposes Code Eval-Instruct that employs various heuristics to increase the complexity of seed code instructions. Magicoder (Wei et al., 2024) proposes to generate new coding problems by drawing inspiration from random snippets collected from open source code.
@@
While previous work has shown significant improvements for models trained on data generated by larger, more competent models, training an LLM on its own generated data is not helpful and can even degrade performance (Zhou et al., 2024). Therefore, to prevent the model from learning errors present in its own generated data, some post-processing steps are essential. For example, Llama 3.1 (Dubey et al., 2024) utilizes error feedback from test execution and adopts an iterative self-correction procedure to revise potential errors. CodeDPO (Zhang et al., 2024) replaces teacher models with a self-generation-and-validation process that uses a PageRank-like algorithm to rank code snippets by correctness and efficiency, yielding diverse preference optimization data without external resources.
@@
In this work, we propose leveraging both positive and negative examples generated by the model, treating pairs of passing and failing responses as chosen-rejected pairs for Direct Preference Optimization (DPO) (Rafailov et al., 2024). Note that our method is complementary to self-correction, rather than orthogonal. By improving the quality of unit tests, our framework enhances the accuracy of unit test execution feedback, and thereby can benefit self-correction scenarios as well. While CodeDPO also employs a self-verification method, it is fundamentally constrained by the quality of its generated tests. In contrast, our approach simultaneously improves the generation of both code and tests in a self-play method, mitigating this bottleneck and enabling more robust training.
@@ 4.3 Evaluating the Base Model
Dataset Acc%${}_{\text{Iter1}}$ Acc%${}_{\text{Iter2}}$ Agreemnt Acc%${}_{\text{Ens.}}$
MBPP 49.50% 51.01% 75.14% 51.12%
LiveCodeBench 37.09% 40.65% 72.38% 40.68%
@@ 4.4 Evaluating Agreement between Iterations
Table 4: Case study of iterative improvement made by Sol-Ver’s test generator. Red highlighted tests refer to incorrect tests.
@@ 6 Conclusion
In this work, we introduced Sol-Ver, a novel self-play solver-verifier framework that iteratively enhances both code and unit test generation capabilities of large language models (LLMs). By enabling the model to act simultaneously as a solver and a verifier, our approach effectively leverages synthetic data to bridge the critical gap between code generation and unit test generation. Experimental results on the Llama 3.1 8B model demonstrate significant improvements in generating both high-quality code solutions and corresponding unit tests without the necessity of human-annotated data or reliance on larger teacher models. Sol-Ver not only advances the state-of-the-art in automated code and test generation but also offers a scalable solution that can be adapted to various programming tasks.

### v3

@@ 1 Introduction
Large language models (LLMs) have demonstrated impressive ability in code generation, significantly enhancing the programming efficiency and productivity of human developers (Li et al., 2022; Roziere et al., 2023; Chaudhary, 2023). The ability to code is largely due to high-quality online coding resources, e.g., coding problem and human-rewritten solutions. However, as these supervised data sources saturate, LLM improvement is diminishing, with data scarcity becoming a key bottleneck for further progress.
@@
To address scarce supervised data for code generation, recent studies use synthetic data techniques like Self-Instruct (Wang et al., 2023) to augmenting LLM training sets. Typically, a high-capacity teacher LLM generates code responses to designed instructions, and this data is then employed to fine-tune a student LLM, thereby enhancing its code generation abilities. Although synthetic code data produced in this manner has demonstrated success, it relies on the availability of a strong teacher model, presumably with a larger parameter size and higher computation costs. Additionally, existing work has shown that training a model on data generated by itself is ineffective because errors introduced during generation tend to accumulate over iterations (Dubey et al., 2024). As a result, there is a critical need for effective methods to verify the generated data.
While evaluating generated code correctness is challenging, often requiring expert human intervention, recent LLM-as-a-judge efforts involve models executing generated code against self-generated unit tests (McAleese et al., 2024; Alshahwan et al., 2024; Dong et al., 2024; Chen et al., 2024, 2022, 2023; Dubey et al., 2024). However, a critical bottleneck emerges: an LLM’s proficiency as a verifier (generating effective unit tests) significantly lags its capability as a solver (generating code solutions), a disparity we quantify in Section 4.3). This gap is largely attributable to the scarcity of high-quality, diverse unit test generation data used during LLM fine-tuning, which predominantly focuses only on code generation.
@@
To address this imbalance and unlock a new avenue for data generation, we introduce Sol-Ver, a self-play solver-verifier framework to iteratively train a model for both code and test generation. The main idea is to let the LLM-as-a-solver and LLM-as-a-verifier help each other. Specifically, we ask the model to generate code solutions and unit tests for the same set of coding problems. By executing the generated test against the generated code, we obtain feedback for training, involving two steps: (1) SFT training: we take the passed examples for fine-tuning the model, and (2) DPO training: we take both passed and failed examples as preference pairs to further train the model aligning with the preference. These training steps are for both code generation and unit test generation, and they can be repeated in an iterative manner.
@@
  • Identification of a critical gap: We empirically demonstrate and analyze the significant gap in LLMs’ abilities between code generation and unit test generation, motivating the need for targeted improvements in test generation.
@@ 2 Related Work
Previous studies applying scaling laws to the training of foundation models have highlighted the critical role of the data size (Kaplan et al., 2020; Hoffmann et al., 2022; Chung et al., 2022). To address the need for larger datasets, synthetic data generation has become a popular and cost-effective solution, which leverages advanced LLMs to produce high-quality data. One notable method is Self-Instruct (Wang et al., 2023; Taori et al., 2023), which employs a pre-trained LLM to generate instruction-output pairs from a small seed dataset.
@@
For code generation, previous work devises synthetic coding instructions using a stronger teacher model (e.g., ChatGPT or GPT-4) and then finetunes a weaker student model (e.g., CodeAlpaca (Chaudhary, 2023) and CodeLlama (Roziere et al., 2023)) with the generated data to distill knowledge from the teacher. For example, code alpaca consists of 20K automatically generated code instructions by applying Self-Instruct on ChatGPT using 21 seed tasks. To enhance the code abilities of LLMs, Luo et al. (2023) proposes Code Eval-Instruct that employs heuristics to increase the complexity of seed code instructions. Magicoder (Wei et al., 2024) proposes to generate coding problems by drawing inspiration from snippets collected from open source code.
@@
While previous work has shown significant improvements for models trained on data generated by larger, more competent models, training an LLM on its own generated data is not helpful and can even degrade performance (Zhou et al., 2024; Alemohammad et al., 2023). To prevent the model from learning errors present in its own generated data, some post-processing steps are essential. For example, CodeT (Chen et al., 2022) leverages pre-trained language models to automatically generate test cases for multiple code samples. It executes the generated code against these test cases and employs a dual execution agreement strategy, ensuring both output consistency and agreement among multiple code solutions. Self-Debug (Chen et al., 2023) enables models to autonomously identify and fix bugs in their generated code. Llama 3.1 (Dubey et al., 2024) utilizes error feedback from execution and adopts an iterative self-correction procedure to revise potential errors. Reinforcement learning from unit test feedback (RLTF) has also been explored to improve code generation by directly optimizing for test pass rates (Le et al., 2022; Ni et al., 2023) CodeDPO (Zhang et al., 2024) replaces teacher models with a self-generation-and-validation process that uses a PageRank-like algorithm to rank code snippets by correctness and efficiency, yielding diverse preference optimization data without external resources. Another related line of work, AutoIF (Li et al., 2024), proposes an iterative framework for generating instruction-following data for general tasks by having an LLM play the roles of instructor and respondent, using another LLM as an evaluator, though it does not specifically focus on code generation or the co-evolution of solver and verifier capabilities.
@@
In this work, we propose leveraging both positive and negative examples generated by the model, treating pairs of passing and failing responses as chosen-rejected pairs for Direct Preference Optimization (DPO) (Rafailov et al., 2024). Note that our method is complementary to self-correction and RLTF, rather than orthogonal. By improving the quality of unit tests, our framework enhances the accuracy of unit test execution feedback, and thereby can benefit self-correction and RLTF scenarios as well.
Sol-Ver Compared to CodeDPO: While CodeDPO (Zhang et al., 2024) also utilizes self-verification for preference data, its efficacy is inherently limited by the initial, and potentially static, quality of the self-generated tests used for verification. If the verifier component is weak, it cannot reliably distinguish high-quality code, potentially leading to suboptimal preference learning. In contrast, Sol-Ver’s self-play mechanism is designed to explicitly address this: it simultaneously improves both the code solver and the test verifier. By iteratively enhancing the verifier’s ability to generate more discerning and comprehensive tests, Sol-Ver breaks this quality ceiling, thereby mitigating the verifier bottleneck and enabling more robust and effective training for both capabilities.
@@ 4.4 Evaluating Agreement between Iterations
Dataset Acc%${}_{\text{Iter1}}$ Acc%${}_{\text{Iter2}}$ Agreemnt Acc%${}_{\text{Ens.}}$
MBPP 49.50% 51.01% 75.14% 51.12%
LiveCodeBench 37.09% 40.65% 72.38% 40.68%
@@
Table 4: Illustrative examples of test case refinement by Sol-Ver’s verifier component across iterations for two distinct programming problems. Red highlighting indicates tests that were incorrect or suboptimal in earlier iterations but were corrected or improved by later iterations, demonstrating the verifier’s learning progress.
@@ 6 Conclusion
In this work, we introduced Sol-Ver, a novel self-play framework designed to address the critical challenge of data scarcity and the performance disparity between LLM-based code solvers and test verifiers. By enabling an LLM to simultaneously embody both roles, Sol-Ver facilitates a co-evolutionary process where improvements in test generation lead to better code synthesis, and vice-versa. This iterative refinement loop generates high-quality synthetic code-test pairs, demonstrably enhancing both capabilities without reliance on human annotations or larger teacher models. Our experiments with Llama 3.1 8B validate Sol-Ver’s efficacy, achieving significant performance gains on established benchmarks. Sol-Ver represents a step towards more autonomous, data-efficient, and robust systems for automated code and test generation, offering a scalable and adaptable paradigm for continuous self-improvement in code AI.
@@
Acknowledgment
We thank Ilia Kulikov for assisting with the coding and answering questions regarding the usage of fairseq2.